### v1

# Advancing Reference-free Evaluation of Video Captions with Factual Analysis

Abstract
Video captions offer concise snapshots of actors, objects, and actions within a video, serving as valuable assets for applications such as question answering and event localization. However, acquiring human annotations for video captions is costly or even impractical, especially when dealing with diverse video domains. Existing models trained on supervised datasets face challenges in evaluating performance across different domains due to the reliance on reference-based evaluation protocols, which necessitate ground truth captions. This assumption is unrealistic for evaluating videos in the wild. To address these limitations, we propose a reference-free evaluation framework that does not require ground truth captions, focusing on factual grounding to ensure accurate assessment of caption quality. We introduce VC-Inspector, a novel caption quality evaluator that is both reference-free and factually grounded. Utilizing large language models, we generate pseudo captions of varying quality based on supervised data, which are subsequently used to train a multimodal model (i.e., Qwen2.5-VL) as the evaluator. Our approach demonstrates superior alignment with human judgments on the VATEX-Eval dataset, outperforming existing methods. The performance also generalizes to image caption datasets, Flickr8K-Expert and Flickr8K-CF, when viewing images as 1-frame videos. Overall, VC-Inspector offers a scalable and generalizable solution for evaluating the factual accuracy of video captions, paving the way for more effective and objective assessment methodologies in diverse video domains.
1 Introduction
Figure 1: Existing reference-free metrics like EMScore [31] often fail to detect factual inaccuracies and lack a consistent scoring scale. VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations with explanations.
Video captioning plays a pivotal role in bridging visual content with natural language, offering concise descriptions of salient entities, actions, and interactions within a video. These captions are instrumental in enabling a wide range of downstream applications, including video-based question answering [27], event localization [28, 12], and content retrieval [8]. To ensure the quality and reliability of captions, robust evaluation metrics are vital for advancing the field of video-language research. However, current captioning models, predominantly trained on supervised datasets [14, 41], rely heavily on reference-based evaluation protocols [26, 3, 19, 33, 39] that compare generated captions against human-written ground truth. While effective in controlled settings, these metrics face significant limitations in real-world scenarios. First, generating such references is prohibitively labor-intensive and costly, often becoming unavailable or infeasible to obtain when scaling across diverse video domains with varying visual and contextual characteristics. Second, these metrics demonstrate a limited ability to understand the semantic meanings of the captions and often fail to recognize valid alternatives.
The challenges of assessing open-domain or in-the-wild videos necessitate the development of evaluation protocols that do not rely on reference captions. However, this remains underexplored within the video research community. Current reference-free metrics [31, 30] typically measure the visual-language semantic alignment using pretrained multimodal embeddings [29], with the comprehension of caption content restricted by the context length of the underlying text encoder. In addition, these metrics frequently overlook factual inaccuracies in the candidate captions, as shown in Figure 1. Furthermore, they are primarily image-based solutions, making them suboptimal for video content.
To address these limitations, we propose a novel reference-free evaluation framework, VC-Inspector, that eliminates the dependency on human-annotated captions. Our approach emphasizes factual grounding, ensuring that the evaluation of caption quality is based on the alignment between the caption and the actual video content, rather than comparison to a predefined reference. We build VC-Inspector atop a lightweight large multimodal model (LMM), equipping the model with the ability to judge the quality of video captions based on factual accuracy, through instruction tuning. However, while supervised video caption datasets are already scarce, captions of imperfect quality are even harder to obtain. We tackle this challenge by introducing a novel data generation pipeline powered by a large language model (LLM), creating pseudo captions of varying quality by altering the factual elements (i.e., objects and actions) in the ground truth captions from ActivityNet [14]. This assembles ActivityNet-FG-It, comprising 44K data points for instruction tuning. Figure 2 illustrates the data generation and training pipeline of VC-Inspector. Unlike previous metrics that offer only a single quality score, our approach also requires the model to explain the rationale behind its scoring, making the evaluator more interpretable.
We start with evaluating VC-Inspector on the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, where the labels are synthetically generated with the same data generation pipeline as the instruction tuning dataset. Results demonstrate that the quality estimates are consistent across datasets, and VC-Inspector is versatile for different visual content and caption lengths. We further evaluate the correlation between the quality scores given by VC-Inspector and human judgments using the VATEX-Eval dataset [31], a standard benchmark for video captioning metric evaluation. The proposed evaluator stands out compared to the previous reference-free metrics, even surpassing most of the reference-based metrics. Finally, we extend the evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], showing the generalization of the proposed metric.
Overall, this work makes the following contributions:
  • We present one training dataset and two evaluation datasets consisting of ground truth captions, synthetic candidate captions, and pseudo quality scores, generated using a novel data generation pipeline.
  • We propose VC-Inspector, an open-source reference-free evaluator featuring factual grounding, with a text context length of 32K tokens and support for long videos.
  • Our evaluator outperforms existing reference-free metrics for video captions, achieving higher correlations with human evaluations and generalizing across visual domains.
2 Related Works
Traditional text-based evaluation can be extended to video captioning by discarding the video and treating the reference caption as ground truth. Alternatively, image-caption metrics can be adapted by sampling images from the video and aggregating image-caption scores. We review these metrics and their limitations below:
Text-only metrics based on references.
Several rule-based metrics have been developed to evaluate generated texts, including widely used methods such as BLEU [26], METEOR [3], ROUGE [19], and CIDEr [33]. While these metrics perform reasonably well in controlled settings, they primarily focus on syntactic structure and fail to capture the semantic meaning of the generated text. To address this limitation, the SPICE series [1, 18] attempts to embed semantic understanding by parsing both reference and candidate captions into objects, attributes, and their relationships. However, these metrics still struggle to account for semantically similar or identical words expressed differently. This persistent challenge has motivated the development of embedding-based evaluation methods, such as BERTScore [40] and its extended version [37], which consider the intrinsic variance between multiple ground truth captions. More recently, CLAIR [5] explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics. However, these metrics all rely exclusively on reference captions for evaluation and do not incorporate the visual input that the captions are meant to describe.
Image-augmented metrics.
To address the limitation of text-only metrics, various approaches have been explored that cross-reference between visual input and captions. VIFIDEL [24] computes semantic similarity by matching object names extracted from the image with those mentioned in the candidate caption. However, objects are discrete representations limited to fixed categories and cannot capture the motion dynamics that are particularly informative for videos. Visual-language models (VLMs), by contrast, provide continuous representations that offer richer semantic alignment between visual and textual modalities. ViLBERTScore [16] extends BERTScore by leveraging the pretrained ViLBERT [23] models, but still requires comparison to reference captions. UMIC [15] and CLIP-based metrics [9, 13, 7, 30, 34, 38, 17, 31, 21] relax this constraint by employing contrastive learning to directly model the semantic alignment between the images and captions. Although promising for short captions, these embedding-based methods are constrained by the context length limitations of their text encoders. Recently, the strong contextual reasoning and instruction following capability of LLMs have been leveraged for evaluation across diverse tasks [22]. Building on this trend, Maeda et al. [25] employ an LLM to compare the candidate caption to the visual context extracted with a VLM. However, these methods only consider static images and do not model the temporal dynamics of video, resulting in suboptimal performance on video caption evaluation.
Figure 2: (left) We present a data generation pipeline designed to systematically create synthetic video captions with diverse quality scores, along with explanations for the assigned scores. (right) This dataset was subsequently used for instruction tuning the VC-Inspector.
In summary, text-based methods require costly, high-quality reference captions, while image-based extensions are suboptimal for video content. These challenges highlight the growing demand for reference-free video caption evaluation tailored for video inputs. However, it remains an underexplored topic in the community, compared to text-based and image-caption evaluation. EMScore [31] was an early attempt that supports evaluating video captions without a reference. Although it considers both frame-level and video-level embeddings, these embeddings are still obtained from an image-based encoder [29], and the text encoder is limited by a short context length. PAC-S [30] and FactVC [21] augment EMScore with positive and negative data synthesis, respectively. While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation. In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality evaluation. More recently, G-VEval [32] extended the G-Eval [22] approach to video by stitching together three frames from each video. However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos. Moreover, the dependence on proprietary models such as GPT-4o limits its scalability and practicality for widespread use. In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
Ground Truth Candidate Caption Rouge-L SPICE Soft-SPICE
The man is feeding a cat on the sofa in the living room The man is feeding a lion on the sofa in the living room 92.31 75.00 78.57
The girl is dancing in the room The girl is sleeping in the room 85.71 66.67 88.19
Table 1: Text-based metrics rely on ground truth for evaluation and often fail to detect factual errors in objects/actions (bolded).
3 Video Caption Quality Estimation
In this section, we provide a concise overview of existing metrics used for evaluating video captions, identify their shortcomings, and outline our objectives.
3.1 Overview
[…]
3.2 Toward Factually Grounded Reference-free Evaluation
In this paper, we aim to develop a factually grounded and reference-free video caption quality evaluator that can generate scores on an absolute scale, and is versatile for various domains and text lengths. We argue that an evaluation protocol is only reliable if it is factually grounded, meaning that the quality assessment must accurately reflect the correctness of object entities and actions in the caption with respect to the video. For example, the evaluator should assign a lower score to a candidate caption where an object is missing or an action is wrong in the video than a correct one, and an even worse score for a caption that has more errors. However, it remains unclear whether existing metrics can effectively capture these subtle changes in factual elements. To investigate this, we probe existing metrics with candidate captions that contain incorrect objects or/and actions. Table 1 and Figure 1 illustrate such cases for reference-based and reference-free metrics, respectively. These results reveal that current metrics are insensitive to these factual errors. Even when captions are semantically distinct (e.g., “cat" vs. “lion" and “dancing" vs. “sleeping"), high scores are assigned due to substantial syntactic overlap.
Motivated by these observations, we propose VC-Inspector, a novel factually grounded and reference-free model for inspecting the video caption quality. We leverage the recent advents of large multimodal models (LMMs) to handle long-context text reasoning and generalized video feature extraction, with the hypothesis that their established efficacy on visual-language joint reasoning can be applied to this task. While evaluation metrics are preferred to be lightweight, we employ the 3B/7B version of Qwen2.5-VL as a model-based evaluator and equip the model with the factual grounding ability through instruction tuning, where the model receives a video $V$ and a candidate caption $\hat{X}$, and generates an integer score ranging from 1 to 5. We also request the model to provide an explanation for the assigned score in the output, similar to [32], which makes the evaluator more interpretable and can serve as a supervision signal for factual grounding. The next section will cover how we systematically generate captions with incorrect factual elements, the corresponding quality estimates, and the explanations.
4 VC-Inspector: Instruction Tuning for Factually Grounded Evaluation
In video captioning, two primary factors that can degrade caption quality are: (1) the inclusion of objects not present in the video, and (2) the introduction of actions not depicted in the video. Our preliminary study indicates that existing metrics fall short in capturing these factual elements in the caption. To address this, we propose to empower LMMs with this ability through instruction tuning. However, it is nontrivial to acquire annotated samples for training. Supervised video caption datasets are already scarce due to the high cost of labeling, and these datasets generally only include correct captions, lacking those of lower quality. To obtain a considerable number of samples with diverse quality, we propose a systematic way to create a synthetic video caption dataset leveraging the internal knowledge of large language models (LLMs).
Figure 3: Data generation pipeline to create a synthetic dataset for training VC-Inspector. While both “talking” and “holding” were identified as actions, only “holding” was sampled for replacement in the synthetic dataset.
4.1 Data Generation
To create captions with incorrect elements, we employed Llama-3.3-70B-Instruct to alter the objects and actions in a ground truth caption from a supervised video caption dataset. Specifically, we created the ActivityNet-FG-It dataset for instruction tuning utilizing the ActivityNet [14] training set. The data generation pipeline is summarized in Figure 3.
[…]
After generating a pseudo caption, it is critical to assign it a reasonable quality score that is intuitive for humans so that it can better mirror humans’ expectations. To achieve this, one could utilize a visual-language model (VLM) to estimate this score; however, VLMs are not robust enough to produce a factually grounded score in a specified range without calibration, and the scoring quality is confined by their capabilities. Instead, we employ a deterministic scoring mechanism based on factual grounding, described as follows,
[…]
$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$ (1)
Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score will reside within the range of 0 and 1. It also guarantees that a caption with more incorrect objects/actions receives a lower score. We then convert the score to an integer in the 1-5 range for better interpretability during model training.
We repeated the aforementioned process to create 10 pseudo captions per ground truth caption, resulting in a total of 374K pseudo captions derived from 37,396 video-caption pairs. The prompts used to guide the generation process are detailed in the supplementary. The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation from the original caption. However, it also naturally yields a skewed and non-uniform score distribution. To mitigate the potential bias during training, we applied a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories. However, due to computational constraints, training with the full 218K instances would require multiple weeks of runtime. Therefore, we further sampled a 44K (8.8K for each label) caption subset from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It. This subset preserves category balance and maintains the diversity and structure of the original data, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100).
4.2 Training
[…]
where we format the explanation $E$ in text using the information collected in the data creation process, i.e., the list of changed objects and actions, detailed in the supplementary. This can serve as extra supervision for the model to implicitly learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time. During training, the model is optimized with the language modeling loss, like in other instruction tuning work [20].
5 Experiments
5.1 Experimental Settings
Datasets.
We train VC-Inspector with the proposed instruction tuning dataset, ActivityNet-FG-It, which comprises 44K video-caption pairs, along with their quality scores and explanations. To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 4.1 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet [14] test set and YouCook2 [41] validation set. However, these are not truly labeled by humans. To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption evaluation quality. It contains 6 captions with varying levels of quality per video, each rated by three human evaluators on a scale of 1 to 5. Since some videos become unavailable on YouTube, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions. Unless otherwise specified, all experiments were evaluated on the same dataset to ensure a fair comparison. We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], by viewing images as single-frame videos to test the generalization of the proposed metric. The former contains 17K image-caption pairs rated by three human experts on a scale of 1 to 4, while the latter collects binary quality assessments for 48K image-caption pairs from crowd sources.
Baselines.
The baselines are organized into three categories: i) Language-based metrics: The evaluation in this category solely relies on the reference texts without considering the visual input. Representative metrics are BLEU [26], ROUGE [19], METEOR [3], CIDEr [33], BERTScore [39], SPICE [1], SPICE-Factual [18], Soft-SPICE  [18] and CLAIR [6]. ii) Image-augmented metrics: These approaches incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore [9], EMScore [31], PAC-S [30], FactVC [21], and G-VEVAL [32]. They usually support both reference-free and reference-based settings, where the former compares the candidate caption directly to the video without reference captions, while the latter uses both visual and textual references. We report results for both settings for completeness, but our primary focus is on the reference-free setting, which is more practical for real-world, in-the-wild videos. iii) Video-based metrics: These methods employ a video encoder to incorporate full video sequences as references. To the best of our knowledge, no existing metric in the literature falls into this category. Therefore, we adapt CLIPScore [9] to the recent advent of ViCLIP [35] as a stronger baseline, ViCLIPScore. Additionally, we compare VC-Inspector against the base model it builds upon, the vanilla Qwen2.5-VL model, to highlight the benefit of fine-tuning for evaluative reasoning.
Metrics.
In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to human judgments. Following prior work [31, 32], we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
Implementation details.
VC-Inspector is developed for two model sizes, 3B and 7B, initialized from their corresponding Qwen2.5-VL pretrained weights. In all experiments, we train the model on 4 NVIDIA-A100 GPUs with a global batch size of 128 and a learning rate of 1e-4. We set both alpha and rank to 32 for the low-rank adaptation with a dropout rate of 0.05. During inference, we use a temperature of 0.0 for reproducibility.
5.2 Generalization of Quality Estimation
In this section, we evaluate the consistency of quality estimates using synthetic data, and validate the effectiveness of VC-Inspector by measuring its alignment with human judgments on the video caption evaluation dataset, VATEX-Eval. Given the limited availability of video caption datasets, we treat images as single-frame videos and extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF, to assess the generalization.
Evaluation with the synthetic data.
ActivityNet-FG-Eval YouCook2-FG-Eval
Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
EMScore [31] 28.94 40.77 20.21 29.24
CLIPScore [9] 28.10 39.65 18.00 26.14
Qwen2.5-VL-3B [2] 37.91 47.80 37.16 47.17
VC-Inspector-3B 49.53 62.01 44.29 55.31
Table 2: Correlation scores on the synthetic ActivityNet-FG-Eval [4] and YouCook2-FG-Eval [14] datasets. The best score is bolded.
Table 2 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies. VC-Inspector, specifically finetuned for factual grounding, consistently outperforms baselines in differentiating incorrect captions. Notably, although it is only trained on ActivityNet-FG-It, its quality estimation generalizes effectively across visual domains. This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs that fail to converge.
Evaluation on VATEX-Eval.
Table 5.2 presents our primary results: human correlation scores on the VATEX-EVAL [31] dataset. Baselines are grouped by reference type as outlined in section 5.1, and the three column-sections correspond to the No Reference, 1-Reference, and 9-Reference settings, respectively. Language-based metrics, which rely heavily on textual references, are not applicable to our target No Reference setting. Therefore, we focus our comparisons on multimodal methods that incorporate visual inputs. VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on CLIP embeddings, including CLIPScore, EMScore, FactVC, and PAC-S. When adapting CLIPScore to the recently introduced ViCLIP [35] (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches. Despite ViCLIPScore being a stronger baseline, it still underperforms relatively to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation. We further compare VC-Inspector with G-VEval [32], a recent LMM-based evaluation method. However, G-VEval is based on GPT-4o, a proprietary, paid model with an estimated 200B parameters. In contrast, VC-Inspector is open-source, lightweight, and reproducible, with configurations available at 3B and 7B parameters depending on system requirements. While having a compact size, our 7B model achieves the highest correlation to human evaluations and could potentially enable on-the-fly quality estimation during training, making it a viable reward model in Reinforcement Learning (RL) applications. Although our primary focus is to compare with the reference-free metrics, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
Evaluation on Flickr8K.
No Reference 1-Reference 9-References
Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
[…]
BLEU_1 [26] - - 12.65 16.52 28.70 36.88
BLEU_4 [26] - - 12.44 36.28 22.76 29.56
ROUGE-L [19] - - 12.94 16.89 23.94 31.06
METEOR [3] - - 16.68 21.80 27.64 35.76
CIDEr [33] - - 17.62 23.02 27.92 36.18
BERTScore [39] - - 15.24 19.82 25.05 32.37
SPICE [1] - - 14.80 18.78 27.41 35.40
SPICE-factual [18] - - 13.59 17.04 26.05 35.58
Soft-SPICE [18] - - 21.25 27.61 36.31 46.41
CLAIR [6] - - 36.00 - 34.80 -
[…]
CLIPScore [9] 22.33 29.09 27.39 35.49 35.21 45.28
EMScore [31] 22.88 29.79 28.63 37.05 36.66 47.00
FactVC [21] 22.79 29.69 28.78 37.22 36.18 46.33
PAC-S [30] 25.10 - 32.60 - 31.40 -
G-VEval [32] 39.40 - 44.90 - 48.10 -
[…]
ViCLIPScore [9, 35] 30.92 39.86 - - - -
Qwen2.5-VL-3B [2] 31.29 36.43 - - - -
Qwen2.5-VL-7B [2] 34.70 39.40 - - - -
VC-Inspector-3B 37.99 42.45 - - - -
VC-Inspector-7B 42.58 45.99 - - - -
Table 3: Human correlation scores on the VATEX-EVAL [31] dataset. indicates results reported from [32] due to GPT-4 licensing issue. The best score for each column section is bolded. Please note that this work focuses on the No Reference setting. Our best model has outperformed all other models in this setting, while remaining competitive with metrics that rely on references.
Metric Flickr8K-Expert Flickr8K-CF
[…]
BLEU_1 [26] 32.20 17.90
BLEU_4 [26] 30.60 16.90
ROUGE [19] 32.10 19.90
METEOR [3] 41.50 22.20
CIDEr [33] 43.60 24.60
SPICE [1] 51.70 24.40
BERTScore [39] - 22.80
CLIPScore [9] 52.60 36.40
PAC-S [30] 55.50 37.60
[…]
CLIPScore [9] 51.10 34.40
PAC-S [30] 53.90 36.00
VC-Inspector-3B 59.86 39.00
VC-Inspector-7B 63.43 45.97
Table 4: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF [10] dataset. The overall best scores are bolded and best of each section is underlined. Our model has outperformed even the reference-based methods.
To expand our evaluation beyond the limited availability of video caption datasets, we treat images as an extreme case of short videos with a single frame. Table 5.2 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF. The former requires a more fine-grained differentiation among the captions, with ratings distributed across 4 levels, whereas the latter employs binary judgments. VC-Inspector is instruction-tuned with captions that exhibit varying degrees of factual inaccuracies, enabling nuanced evaluation of caption quality. This allows the model to perform effectively across both benchmarks, despite their differing rating scheme. In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, substantially narrowing the gap to inter-human correlation ($\tau_{b}\approx 73$[1], and even outperforming several reference-based metrics. These results demonstrate the strong generalization capability of VC-Inspector across visual domains and video lengths.
5.3 Ablation and Analysis
We conduct ablation studies on VC-Inspector components on the VATEX-Eval benchmark:
Data synthesis strategies.
Since both objects and actions are informative factual elements for a video, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components. To this end, during the data generation process (outlined in section 4.1), we systematically altered both elements in ground truth captions to create pseudo captions for training. In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both. As shown in Table 5, all variants achieve strong alignment with human ratings compared to EMScore. However, the variant that alters both objects and actions yields the best performance. These results highlight the importance of both factual elements, object and action, in capturing the context of video content for caption evaluation. They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
Role of explanations.
Figure 4: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others). VC-Inspector produces quality assessments highly consistent with ground truth scores, along with explanatory insights into factual errors (highlighted in red).
Metric Data Synthesis $\tau_{b}$ $\rho$
EMScore [31] n/a 22.88 29.79
VC-Inspector-3B Change objects only 36.40 41.20
Change actions only 33.23 39.63
Change both (Ours) 37.99 42.45
Table 5: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval.
During our data generation process, we systematically synthesize pseudo captions along with their corresponding quality estimates. This process yields an informative “side product": explanations that indicate where factual errors are located within the candidate captions. We leverage these explanations as an auxiliary supervision signal during training, enabling the model to implicitly learn factual grounding, which has been shown to be effective compared to the variant without explanations in Table 6. Not only are explanations useful in training, but they also enhance the interpretability of the model-based metric, providing transparency of why the scores are assigned to given captions. Further evaluation of the explanation quality is provided in the supplementary.
Metric stability.
To evaluate the stability of our metric, following prior work [36], we computed the Pearson correlation between scores obtained across two runs. We find a perfect correlation of 1.0 (as expected with temperature 0.0), confirming that VC-Inspector produces stable quality estimation across multiple runs.
VC-Inspector-3B $\tau_{b}$ $\rho$
Without Explanations 34.29 38.18
With Explanations 37.99 42.45
Table 6: Impact of explanations on the model performance. In the “Without Explanation" setting, we have trained the model only on the score, removing the pseudo explanations.
Computational efficiency.
We assess the computational efficiency by comparing the average runtime per video clip. EMScore [31], ViCLIPScore [35], and VC-Inspector require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU. VC-Inspector demonstrates superior efficiency compared to existing methods.
5.4 Qualitative Results
Figure 4 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (middle, bottom) datasets. Without relying on reference captions, the model evaluates candidate captions based on the associated video content and produces quality scores that closely mirror human judgments. The explanation pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns factually grounded scores accordingly. Additional qualitative results are provided in the supplementary.
6 Conclusion and Discussion
This work addresses the challenge of evaluating video captions across diverse domains without relying on human-annotated reference captions. We identified the limitations of existing metrics, and proposed a novel reference-free and factually grounded evaluation framework, VC-Inspector, based on LMMs. By delving into the factual analysis of captions, we empower the LMM with the factual grounding ability through instruction tuning, where we systematically create pseudo captions of diverse quality using an LLM. Experimental results across multiple domains demonstrate that VC-Inspector achieves high consistency with human evaluators, and outperforms existing metrics on detecting factual errors. Its versatility and interpretability make it a practical tool for evaluating the factual accuracy of video captions in real-world settings.
While our current focus is on factual errors, particularly object and action substitutions, the proposed paradigm can be potentially extended to simulate other types of errors (e.g., attribute replacements), as supported by our ablations on synthesizing only one type of error. Additionally, we recognize the importance of evaluating temporal coherence and narrative structure, which remain underexplored. Future work should expand this paradigm to simulate and detect a broader range of errors, enabling more comprehensive and context-aware evaluation strategies for video captions.
Appendix
The appendix is organized as follows. Section A lists the detailed prompts used in the proposed data generation pipeline, model fine-tuning, and explanation evaluation. Section B provides the evaluation of explanations generated by VC-Inspector, and section C presents additional visual examples with candidate captions of diverse quality.
Appendix A Prompts
The data generation prompts are reported on the following blocks:
  • Extract object – Prompt A
  • Extract action – Prompt A
  • Find similar object – Prompt A
  • Find similar action – Prompt A
  • Substitute object or action – Prompt A
The fine-tuning prompt and the prompt for evaluating the generated explanations are reported in Prompt A and Prompt A, respectively.
Prompt A.1: Prompt to extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", …]. Don’t include any verbs. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.2: Prompt to extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", …]. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.3: Prompt to find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent. The new class cannot be a synonym or similar terms to the original object. It can be an antonym or any co-hyponym. For example, generate "dog" for "cat". ONLY REPLY THE NEW CLASS.
### Input: {{object}}
### Output:
Prompt A.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action. The new action cannot be a synonym or similar terms to the original action. For example, generate "put into" for "take out of". ONLY REPLY THE NEW ACTION.
### Input: {{action}}
### Output:
Prompt A.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act}} in {{caption}} as {{new_obj_act}} . Keep the answer in the same format as {{caption}} . ONLY REPLY THE ANSWER.
### Input: cap
### Output:
Prompt A.6: Fine-tuning prompt ### USER:
[[VIDEO]]
<caption> {{caption}} </caption> You are given a video and a caption describing the video content. Please rate the helpfulness, relevance, accuracy, level of details of the caption. The overall score should be on a scale of 1 to 5, where a higher score indicates better overall performance. Please first output a single line containing only one integer indicating the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias. STRICTLY FOLLOW THE FORMAT.
### ASSISTANT:
{{quality_score}}
The caption does not accurately capture the video content. For example, the actions ( {{wrong_act}} ) are incorrect / the objects ( {{wrong_obj}} ) are incorrect/the objects ( {{wrong_obj}} ) and actions ( {{wrong_act}} ) are incorrect.
Prompt A.7: Prompt for explanation evaluation [Context]
{{ground_truth_caption}}
[Caption]
{{caption_to_evaluate}}
[Groundtruth]
{{ground_truth_explanation}}
[End of Groundtruth]
[Assistant]
{{predicted_explanation}}
[End of Assistant]
[System]
We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video. For your reference, the visual content in the video is represented with a few sentences describing the same video. You are also given a ground truth evaluation to that caption. Please rate the helpfulness, relevance, accuracy, level of details of the response by comparing to the ground truth and referring to the context information. Provide an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
Appendix B Evaluation on Explanation Quality
We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations. Table 7 presents two metrics for this experiment. The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings. On the other hand, the LLM Score is derived by using the prompt detailed in section A to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to [20]. We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations. Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
Dataset BERT Score LLM Score
ActivityNet-FG-Eval 0.79 93.11
YouCook2-FG-Eval 0.70 90.97
Table 7: Evaluation of the explanations generated by VC-Inspector-3B on two synthetic evaluation datasets.
Figure 5: Additional visual examples on VATEX-Eval. VC-Inspector identifies the incorrect objects (highlighted in red) and assigns scores aligned with human evaluators.
Figure 6: Additional visualization of VC-Inspector results on ActivityNet-FG-Eval videos, with candidate captions of diverse quality. Incorrect objects and actions are identified by VC-Inspector and labeled in red.
Appendix C Additional Visual Examples
Figure 5 presents additional visual examples of VC-Inspector applied to VATEX-Eval. Consistent with the results shown in Figure 4 of the main paper, VC-Inspector effectively identifies incorrect objects and actions in candidate captions, assigning quality scores that closely align with human judgments. In Figure 6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual inaccuracies. The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual errors.

### v2

@@
-Advancing Reference-free Evaluation of Video Captions with Factual Analysis
+VC-Inspector: Advancing Reference-free Evaluation of Video Captions with Factual Analysis
 Abstract
-Video captions offer concise snapshots of actors, objects, and actions within a video, serving as valuable assets for applications such as question answering and event localization.
-However, acquiring human annotations for video captions is costly or even impractical, especially when dealing with diverse video domains.
-Existing models trained on supervised datasets face challenges in evaluating performance across different domains due to the reliance on reference-based evaluation protocols, which necessitate ground truth captions.
-This assumption is unrealistic for evaluating videos in the wild.
-To address these limitations, we propose a reference-free evaluation framework that does not require ground truth captions, focusing on factual grounding to ensure accurate assessment of caption quality.
-We introduce VC-Inspector, a novel caption quality evaluator that is both reference-free and factually grounded.
-Utilizing large language models, we generate pseudo captions of varying quality based on supervised data, which are subsequently used to train a multimodal model (i.e., Qwen2.5-VL) as the evaluator.
-Our approach demonstrates superior alignment with human judgments on the VATEX-Eval dataset, outperforming existing methods.
-The performance also generalizes to image caption datasets, Flickr8K-Expert and Flickr8K-CF, when viewing images as 1-frame videos.
-Overall, VC-Inspector offers a scalable and generalizable solution for evaluating the factual accuracy of video captions, paving the way for more effective and objective assessment methodologies in diverse video domains.
+We propose VC-Inspector, a lightweight, open-source large multimodal model (LMM) for reference-free evaluation of video captions, with a focus on factual accuracy.
+Unlike existing metrics that suffer from limited context handling, weak factuality assessment, or reliance on proprietary services, VC-Inspector offers a reproducible, fact-aware alternative that aligns closely with human judgments.
+To enable robust training and interpretable evaluation, we introduce a systematic approach for generating captions with controllable errors, paired with graded quality scores and explanatory annotations.
+Experiments show that VC-Inspector achieves state-of-the-art correlation with human judgments, generalizing across diverse domains (e.g., VATEX-Eval, Flickr8K-Expert, and Flickr8K-CF benchmarks) and revealing the potential for caption improvement.
+[^1] https://dipta007.github.io/VC-Inspector/
+VC-Inspector: Advancing Reference-free Evaluation
+of Video Captions with Factual Analysis
+Shubhashis Roy Dipta  1thanks: These authors contributed equally to this work.,  Tz-Ying Wu  2∗,  Subarna Tripathi  2 1University of Maryland, Baltimore County2Intel sroydip1@umbc.edu
 1 Introduction
-Figure 1: Existing reference-free metrics like EMScore [31] often fail to detect factual inaccuracies and lack a consistent scoring scale.
-VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations with explanations.
-Video captioning plays a pivotal role in bridging visual content with natural language, offering concise descriptions of salient entities, actions, and interactions within a video.
-These captions are instrumental in enabling a wide range of downstream applications, including video-based question answering [27], event localization [28, 12], and content retrieval [8].
-To ensure the quality and reliability of captions, robust evaluation metrics are vital for advancing the field of video-language research.
-However, current captioning models, predominantly trained on supervised datasets [14, 41], rely heavily on reference-based evaluation protocols [26, 3, 19, 33, 39] that compare generated captions against human-written ground truth.
-While effective in controlled settings, these metrics face significant limitations in real-world scenarios.
-First, generating such references is prohibitively labor-intensive and costly, often becoming unavailable or infeasible to obtain when scaling across diverse video domains with varying visual and contextual characteristics.
-Second, these metrics demonstrate a limited ability to understand the semantic meanings of the captions and often fail to recognize valid alternatives.
-The challenges of assessing open-domain or in-the-wild videos necessitate the development of evaluation protocols that do not rely on reference captions.
-However, this remains underexplored within the video research community.
-Current reference-free metrics [31, 30] typically measure the visual-language semantic alignment using pretrained multimodal embeddings [29], with the comprehension of caption content restricted by the context length of the underlying text encoder.
-In addition, these metrics frequently overlook factual inaccuracies in the candidate captions, as shown in Figure 1.
-Furthermore, they are primarily image-based solutions, making them suboptimal for video content.
-To address these limitations, we propose a novel reference-free evaluation framework, VC-Inspector, that eliminates the dependency on human-annotated captions.
-Our approach emphasizes factual grounding, ensuring that the evaluation of caption quality is based on the alignment between the caption and the actual video content, rather than comparison to a predefined reference.
-We build VC-Inspector atop a lightweight large multimodal model (LMM), equipping the model with the ability to judge the quality of video captions based on factual accuracy, through instruction tuning.
-However, while supervised video caption datasets are already scarce, captions of imperfect quality are even harder to obtain.
-We tackle this challenge by introducing a novel data generation pipeline powered by a large language model (LLM), creating pseudo captions of varying quality by altering the factual elements (i.e., objects and actions) in the ground truth captions from ActivityNet [14].
-This assembles ActivityNet-FG-It, comprising 44K data points for instruction tuning.
-Figure 2 illustrates the data generation and training pipeline of VC-Inspector.
-Unlike previous metrics that offer only a single quality score, our approach also requires the model to explain the rationale behind its scoring, making the evaluator more interpretable.
-We start with evaluating VC-Inspector on the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, where the labels are synthetically generated with the same data generation pipeline as the instruction tuning dataset.
-Results demonstrate that the quality estimates are consistent across datasets, and VC-Inspector is versatile for different visual content and caption lengths.
-We further evaluate the correlation between the quality scores given by VC-Inspector and human judgments using the VATEX-Eval dataset [31], a standard benchmark for video captioning metric evaluation.
-The proposed evaluator stands out compared to the previous reference-free metrics, even surpassing most of the reference-based metrics.
-Finally, we extend the evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], showing the generalization of the proposed metric.
-Overall, this work makes the following contributions:
-  • We present one training dataset and two evaluation datasets consisting of ground truth captions, synthetic candidate captions, and pseudo quality scores, generated using a novel data generation pipeline.
-  • We propose VC-Inspector, an open-source reference-free evaluator featuring factual grounding, with a text context length of 32K tokens and support for long videos.
-  • Our evaluator outperforms existing reference-free metrics for video captions, achieving higher correlations with human evaluations and generalizing across visual domains.
+Video captions summarize salient objects and actions in videos, supporting downstream applications such as question answering, event localization, and retrieval P.J.
+and Kovoor (2024); Qian et al.
+(2024); Gabeur et al.
+(2020).
+Most evaluation metrics rely on reference-based protocols Papineni et al.
+(2002); Banerjee and Lavie (2005); Lin (2004); Vedantam et al.
+(2015); Zhang et al.
+(2020), which compare captions to human-written references.
+These approaches are costly to scale and often fail to capture semantic equivalence.
+Figure 1: Existing reference-free metrics like EMScore Shi et al.
+(2022) often fail to detect factual inaccuracies and lack a consistent scoring scale.
+VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations.
+Evaluating captions for in-the-wild videos requires reference-free protocols that do not rely on ground truth captions, yet this remains underexplored.
+Existing metrics Shi et al.
+(2022); Sarto et al.
+(2023) typically measure visual-language alignment using pretrained multimodal embeddings Radford et al.
+(2021), but are limited by text encoder context length, and lack a consistent scoring scale, making interpretation difficult ( Fig.˜1).
+Recently, large proprietary models such as GPT-4o are used to rate captions on a fixed scale Tong et al.
+(2025), but these approaches depend heavily on prompt engineering and are not fully reproducible.
+Additionally, most prior methods are image-centric, rendering them suboptimal for video content.
+In this work, we aim to develop a reference-free evaluation metric for video captions that removes dependency on human-annotated references while remaining robust and intuitive for humans.
+Our approach is grounded in factual accuracy, as factual elements like objects and actions are critical for video understanding.
+Ideally, a reliable metric should degrade scores proportionally to factual errors.
+For example, for a video showing a little girl sitting on a chair, the caption “A little girl is sleeping on a chair” should receive a higher score than “A woman is sleeping on a chair”, because the latter deviates more from the video content.
+However, existing metrics like EMScore Shi et al.
+(2022) often fail to capture even basic factual inaccuracies, such as incorrect objects/actions.
+To address these limitations, we propose VC-Inspector, built on top of a lightweight, open-source large multimodal model (LMM) trained to assess caption quality based on factual correctness.
+Unlike previous metrics that provide only a score, VC-Inspector also generates explanations for its judgments, improving interpretability.
+One key challenge is the lack of captions with varying degrees of factual quality for training.
+We overcome this by introducing a novel data generation framework powered by a large language model (LLM), which controllably modifies factual elements in ground-truth captions from ActivityNet Krishna et al.
+(2017).
+This process yields ActivityNet-FG-It, a dataset of 44K examples for instruction tuning.
+Fig.˜2 illustrates our data generation and training framework.
+We first evaluate VC-Inspector on two synthetic datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, generated using the same pipeline as our instruction-tuning dataset.
+Results show that VC-Inspector produces consistent quality estimates across datasets, showing versatility for diverse visual content and caption lengths.
+Next, we assess correlation with human judgments on VATEX-Eval Shi et al.
+(2022), a standard benchmark for video captioning metric evaluation.
+VC-Inspector outperforms prior reference-free metrics and even surpasses most reference-based metrics.
+Finally, we extend evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF Hodosh et al.
+(2013), highlighting the strong generalization capability of our approach.
+Beyond the evaluation of the generated scores, we further show that the explanatory outputs of VC-Inspector not only enhance quality estimation but also contribute to caption refinement.
+To summarize, this work makes the following contributions:
+  • We introduce a scalable pipeline for synthesizing video captions with controllable factual errors, enabling large-scale training and evaluation, without costly human annotation.
+  • We propose VC-Inspector, a fact-aware, reference-free video caption evaluator that jointly predicts quality scores and factual error explanations, enhancing interpretability and guiding caption improvement.
+  • We demonstrate that VC-Inspector achieves high correlations with human judgments, outperforming existing reference-free methods and rivaling reference-based metrics across video and image caption benchmarks, showing strong cross-domain generalization.
 2 Related Works
-Traditional text-based evaluation can be extended to video captioning by discarding the video and treating the reference caption as ground truth.
-Alternatively, image-caption metrics can be adapted by sampling images from the video and aggregating image-caption scores.
-We review these metrics and their limitations below:
 Text-only metrics based on references.
-Several rule-based metrics have been developed to evaluate generated texts, including widely used methods such as BLEU [26], METEOR [3], ROUGE [19], and CIDEr [33].
-While these metrics perform reasonably well in controlled settings, they primarily focus on syntactic structure and fail to capture the semantic meaning of the generated text.
-To address this limitation, the SPICE series [1, 18] attempts to embed semantic understanding by parsing both reference and candidate captions into objects, attributes, and their relationships.
-However, these metrics still struggle to account for semantically similar or identical words expressed differently.
-This persistent challenge has motivated the development of embedding-based evaluation methods, such as BERTScore [40] and its extended version [37], which consider the intrinsic variance between multiple ground truth captions.
-More recently, CLAIR [5] explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics.
-However, these metrics all rely exclusively on reference captions for evaluation and do not incorporate the visual input that the captions are meant to describe.
+Traditional rule-based metrics such as BLEU Papineni et al.
+(2002), METEOR Banerjee and Lavie (2005), ROUGE Lin (2004), and CIDEr Vedantam et al.
+(2015) primarily capture syntactic similarity and fail to reflect semantic meaning.
+SPICE Anderson et al.
+(2016); Li et al.
+(2023) introduces semantic parsing of objects, attributes, and relations for both reference and candidate captions, but struggle with paraphrases.
+Embedding-based methods, such as BERTScore Zhang et al.
+(2020) and its variant Yi et al.
+(2020), which consider the intrinsic variance between multiple ground truth captions.
+More recently, CLAIR Chan et al.
+(2023a) explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics.
+However, all these metrics depend on reference captions for evaluation and ignore the visual content.
 Image-augmented metrics.
-To address the limitation of text-only metrics, various approaches have been explored that cross-reference between visual input and captions.
-VIFIDEL [24] computes semantic similarity by matching object names extracted from the image with those mentioned in the candidate caption.
-However, objects are discrete representations limited to fixed categories and cannot capture the motion dynamics that are particularly informative for videos.
-Visual-language models (VLMs), by contrast, provide continuous representations that offer richer semantic alignment between visual and textual modalities.
-ViLBERTScore [16] extends BERTScore by leveraging the pretrained ViLBERT [23] models, but still requires comparison to reference captions.
-UMIC [15] and CLIP-based metrics [9, 13, 7, 30, 34, 38, 17, 31, 21] relax this constraint by employing contrastive learning to directly model the semantic alignment between the images and captions.
-Although promising for short captions, these embedding-based methods are constrained by the context length limitations of their text encoders.
-Recently, the strong contextual reasoning and instruction following capability of LLMs have been leveraged for evaluation across diverse tasks [22].
-Building on this trend, Maeda et al.
-[25] employ an LLM to compare the candidate caption to the visual context extracted with a VLM.
-However, these methods only consider static images and do not model the temporal dynamics of video, resulting in suboptimal performance on video caption evaluation.
+To overcome the limitation of text-only metrics, image-aware approaches align captions with visual input.
+VIFIDEL Madhyastha et al.
+(2019) matches object names between images and captions, but is restricted to discrete categories and cannot capture motion dynamics in videos.
+Visual-language model (VLM) based metrics offer richer semantic alignment through continuous representations.
+ViLBERTScore Lee et al.
+(2020) extends BERTScore using ViLBERT Lu et al.
+(2019), but still requires comparison to reference captions.
+UMIC Lee et al.
+(2021) and CLIP-based metrics Hessel et al.
+(2022); Jiang et al.
+(2019); Cui et al.
+(2018); Sarto et al.
+(2023); Wada et al.
+(2024); Zeng et al.
+(2024); Lee et al.
+(2024); Shi et al.
+(2022); Liu and Wan (2023) relax this constraint by directly modeling image-caption semantic alignment via contrastive learning, though remain limited by the context length of their text encoders.
+Recently, Maeda et al.
+(2024) compares the candidate caption to the VLM-derived visual context with a LLM.
+However, these methods focus on static images and do not model the temporal dynamics, which limits their effectivenss for video caption evaluation.
+In summary, text-based methods require costly, high-quality reference captions, while image-based extensions remain less effective for video content.
+These challenges underscore the need for reference-free video caption evaluation tailored to video inputs, a topic that is relatively less studied compared to text-based and image-caption evaluation.
+EMScore Shi et al.
+(2022) was an early attempt that supports evaluating video captions without a reference.
+Although it considers both frame-level and video-level embeddings, they are derived from an image-based encoder Radford et al.
+(2021), and its text encoder is constrained by short context length.
+PAC-S Sarto et al.
+(2023) and FactVC Liu and Wan (2023) augment EMScore with positive and negative data synthesis, respectively.
+While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation.
+In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality assessment.
+More recently, G-VEval Tong et al.
+(2025) extends G-Eval Liu et al.
+(2023b) to video by stitching together three frames per video.
+However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos.
+Moreover, the dependence on large proprietary models such as GPT-4o limits its scalability and reproducibility for widespread use.
+In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
+3 Methodology
 Figure 2: (left) We present a data generation pipeline designed to systematically create synthetic video captions with diverse quality scores, along with explanations for the assigned scores.
 (right) This dataset was subsequently used for instruction tuning the VC-Inspector.
-In summary, text-based methods require costly, high-quality reference captions, while image-based extensions are suboptimal for video content.
-These challenges highlight the growing demand for reference-free video caption evaluation tailored for video inputs.
-However, it remains an underexplored topic in the community, compared to text-based and image-caption evaluation.
-EMScore [31] was an early attempt that supports evaluating video captions without a reference.
-Although it considers both frame-level and video-level embeddings, these embeddings are still obtained from an image-based encoder [29], and the text encoder is limited by a short context length.
-PAC-S [30] and FactVC [21] augment EMScore with positive and negative data synthesis, respectively.
-While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation.
-In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality evaluation.
-More recently, G-VEval [32] extended the G-Eval [22] approach to video by stitching together three frames from each video.
-However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos.
-Moreover, the dependence on proprietary models such as GPT-4o limits its scalability and practicality for widespread use.
-In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
-Ground Truth Candidate Caption Rouge-L SPICE Soft-SPICE
-The man is feeding a cat on the sofa in the living room The man is feeding a lion on the sofa in the living room 92.31 75.00 78.57
-The girl is dancing in the room The girl is sleeping in the room 85.71 66.67 88.19
-Table 1: Text-based metrics rely on ground truth for evaluation and often fail to detect factual errors in objects/actions (bolded).
-3 Video Caption Quality Estimation
-In this section, we provide a concise overview of existing metrics used for evaluating video captions, identify their shortcomings, and outline our objectives.
 3.1 Overview
 […]
-3.2 Toward Factually Grounded Reference-free Evaluation
-In this paper, we aim to develop a factually grounded and reference-free video caption quality evaluator that can generate scores on an absolute scale, and is versatile for various domains and text lengths.
-We argue that an evaluation protocol is only reliable if it is factually grounded, meaning that the quality assessment must accurately reflect the correctness of object entities and actions in the caption with respect to the video.
-For example, the evaluator should assign a lower score to a candidate caption where an object is missing or an action is wrong in the video than a correct one, and an even worse score for a caption that has more errors.
-However, it remains unclear whether existing metrics can effectively capture these subtle changes in factual elements.
-To investigate this, we probe existing metrics with candidate captions that contain incorrect objects or/and actions.
-Table 1 and Figure 1 illustrate such cases for reference-based and reference-free metrics, respectively.
-These results reveal that current metrics are insensitive to these factual errors.
-Even when captions are semantically distinct (e.g., “cat" vs.
-“lion" and “dancing" vs.
-“sleeping"), high scores are assigned due to substantial syntactic overlap.
-Motivated by these observations, we propose VC-Inspector, a novel factually grounded and reference-free model for inspecting the video caption quality.
-We leverage the recent advents of large multimodal models (LMMs) to handle long-context text reasoning and generalized video feature extraction, with the hypothesis that their established efficacy on visual-language joint reasoning can be applied to this task.
-While evaluation metrics are preferred to be lightweight, we employ the 3B/7B version of Qwen2.5-VL as a model-based evaluator and equip the model with the factual grounding ability through instruction tuning, where the model receives a video $V$ and a candidate caption $\hat{X}$, and generates an integer score ranging from 1 to 5.
-We also request the model to provide an explanation for the assigned score in the output, similar to [32], which makes the evaluator more interpretable and can serve as a supervision signal for factual grounding.
-The next section will cover how we systematically generate captions with incorrect factual elements, the corresponding quality estimates, and the explanations.
-4 VC-Inspector: Instruction Tuning for Factually Grounded Evaluation
-In video captioning, two primary factors that can degrade caption quality are: (1) the inclusion of objects not present in the video, and (2) the introduction of actions not depicted in the video.
-Our preliminary study indicates that existing metrics fall short in capturing these factual elements in the caption.
-To address this, we propose to empower LMMs with this ability through instruction tuning.
-However, it is nontrivial to acquire annotated samples for training.
-Supervised video caption datasets are already scarce due to the high cost of labeling, and these datasets generally only include correct captions, lacking those of lower quality.
-To obtain a considerable number of samples with diverse quality, we propose a systematic way to create a synthetic video caption dataset leveraging the internal knowledge of large language models (LLMs).
+[…]
+We argue that a reliable evaluation protocol should be factually grounded, interpretable, and scalable.
+It must reflect the correctness of factual elements (e.g., objects and actions) in the caption relative to the video.
+For example, captions with missing objects or incorrect actions should receive lower scores, and those with multiple errors should be penalized more severely.
+Furthermore, evaluation should be intuitive for humans and reproducible for broad adoption.
+To address these limitations, we propose VC-Inspector, which leverages a lightweight, open-source large multimodal model (LMM) as the backbone for long-context reasoning and generalized video feature extraction.
+Our hypothesis is that the strong visual-language reasoning capability of LMMs can transfer effectively to this task.
+However, pretrained LMMs do not inherently produce factually grounded ratings and require targeted supervision (Table˜2).
+A key challenge is the scarcity of annotated captions exhibiting diverse degrees of factual accuracy for effective training.
+To overcome this, we introduce a novel data generation pipeline powered by large language models (LLMs) that systematically synthesizes candidate captions along with quality estimates and factual error explanations.
+The overall data generation and training framework is illustrated in Fig.˜2.
+3.2 Data Generation
+To create captions with controlled factual errors, we employ Llama-3.3-70B-Instruct to systematically alter objects and actions in ground truth captions from a supervised video caption dataset.
+Using this approach, we construct ActivityNet-FG-It, an instruction-tuning dataset derived from the ActivityNet Krishna et al.
+(2017) training set.
+The data generation pipeline is illustrated in Fig.˜3.
+Caption generation.
+Given a ground truth caption $X$, we first prompt the LLM to extract the set of objects ${\cal O}=\{o_{1},...,o_{M}\}$ and actions ${\cal A}=\{a_{1},...,a_{N}\}$.
+We then randomly sample $K\sim Unif(0,M)$ objects and $L\sim Unif(0,N)$ actions to replace, forming a subset $\cal{R}\subseteq{\cal O}\cup{\cal A}$, where $|{\cal R}|=K+L$ and $Unif(a,b)$ denotes a discrete uniform distribution over integers $\{a,...,b\}$.
+[…]
+This process also yields a factual error explanation for $\tilde{X}$, enabling fine-grained supervision. The detailed prompts and explanation formats are provided in App.˜D.
 Figure 3: Data generation pipeline to create a synthetic dataset for training VC-Inspector.
 While both “talking” and “holding” were identified as actions, only “holding” was sampled for replacement in the synthetic dataset.
-4.1 Data Generation
-To create captions with incorrect elements, we employed Llama-3.3-70B-Instruct to alter the objects and actions in a ground truth caption from a supervised video caption dataset.
-Specifically, we created the ActivityNet-FG-It dataset for instruction tuning utilizing the ActivityNet [14] training set.
-The data generation pipeline is summarized in Figure 3.
+Scoring.
+After generating a pseudo caption, assigning an intuitive quality score is essential to align with human expectations.
+While pretrained embeddings like visual-language model (VLM) can estimate this score, they lack a fixed scale, are constrained by context length, and do not guarantee factual grounding.
+Instead, we adopt a deterministic scoring mechanism based on factual accuracy:
 […]
-After generating a pseudo caption, it is critical to assign it a reasonable quality score that is intuitive for humans so that it can better mirror humans’ expectations.
-To achieve this, one could utilize a visual-language model (VLM) to estimate this score; however, VLMs are not robust enough to produce a factually grounded score in a specified range without calibration, and the scoring quality is confined by their capabilities.
-Instead, we employ a deterministic scoring mechanism based on factual grounding, described as follows,
-[…]
-$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$ (1)
-Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score will reside within the range of 0 and 1.
-It also guarantees that a caption with more incorrect objects/actions receives a lower score.
-We then convert the score to an integer in the 1-5 range for better interpretability during model training.
-We repeated the aforementioned process to create 10 pseudo captions per ground truth caption, resulting in a total of 374K pseudo captions derived from 37,396 video-caption pairs.
-The prompts used to guide the generation process are detailed in the supplementary.
-The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation from the original caption.
-However, it also naturally yields a skewed and non-uniform score distribution.
-To mitigate the potential bias during training, we applied a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories.
-However, due to computational constraints, training with the full 218K instances would require multiple weeks of runtime.
-Therefore, we further sampled a 44K (8.8K for each label) caption subset from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It.
-This subset preserves category balance and maintains the diversity and structure of the original data, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100).
-4.2 Training
+$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$
+Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score lies on a fixed scale between 0 and 1, ensuring captions with more incorrect elements receive lower scores.
+Since LLMs are inherently unreliable when comparing floating-point values (Spathis and Kawsar, 2023), we discretize the score into a 1-5 range ($score=round(score\times 4+1)$), consistent with standard human annotation protocols (e.g., Likert-style ratings in VATEX-Eval).
+While this may induce minor loss of information (e.g., $0.87\rightarrow 4$ and $0.88\rightarrow 5$), this conversion preserves ordinal relationships and error severity while maintaining the underlying deterministic structure.
+Moreover, VC-Inspector is jointly trained to produce textual explanations that explicitly identify incorrect objects and actions, providing fine-grained supervision beyond the scalar score.
+Post-processing.
+We repeat the caption generation and scoring process to create 10 pseudo captions per ground truth caption, yielding 374K pseudo captions derived from 37,396 video-caption pairs.
+The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation.
+However, this naturally results in a skewed and non-uniform score distribution.
+To mitigate the potential bias during training, we apply a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories.
+Due to computational constraints, training on the full 218K instances would require multiple weeks.
+Therefore, we further sample 44K captions (8.8K per label) from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It.
+This subset preserves category balance and maintains the diversity, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100).
+Importantly, although we use Llama as the generator and perform subsampling as a practical design choice, our data generation pipeline is model-agnostic and applies across model scales and caption datasets, enabling scalable dataset construction at arbitrary sizes.
+3.3 Training of VC-Inspector
 […]
 […]
-where we format the explanation $E$ in text using the information collected in the data creation process, i.e., the list of changed objects and actions, detailed in the supplementary.
-This can serve as extra supervision for the model to implicitly learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time.
-During training, the model is optimized with the language modeling loss, like in other instruction tuning work [20].
+The explanation $E$ is formatted in natural language using information collected during data generation (i.e., the list of changed objects and actions), as detailed in App.˜D.
+This can serve as extra supervision for the model to learn factual grounding and provide interpretable reasoning for this model-based evaluator at test time.
+Fig.˜4 presents an example of the VC-Inspector output.
+During training, the model is optimized with the standard language modeling loss Liu et al.
+(2023a).
+Figure 4: Visual example on VATEX-Eval. VC-Inspector produces quality assessments consistent with ground truth scores, and factual error explanations (highlighted in red). More examples are in App.˜A.
 4 Experiments
@@
 We train VC-Inspector with the proposed instruction tuning dataset, ActivityNet-FG-It, which comprises 44K video-caption pairs, along with their quality scores and explanations.
-To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 4.1 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet [14] test set and YouCook2 [41] validation set.
+To evaluate whether VC-Inspector can generalize to other videos and visual domains, we further follow the same data generation pipeline in section 3.2 to create two evaluation datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, according to the ActivityNet Krishna et al.
+(2017) test set and YouCook2 Zhou et al.
+(2018) validation set.
 However, these are not truly labeled by humans.
-To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption evaluation quality.
+To evaluate whether the quality scores given by VC-Inspector are aligned with human evaluators, we utilized the widely adopted VATEX-EVAL dataset, specifically designed for evaluating video caption quality.
 It contains 6 captions with varying levels of quality per video, each rated by three human evaluators on a scale of 1 to 5.
-Since some videos become unavailable on YouTube, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions.
+Since some videos become unavailable, we collect the remaining subset of 2,590 videos and the corresponding 15,540 candidate captions.
 Unless otherwise specified, all experiments were evaluated on the same dataset to ensure a fair comparison.
-We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], by viewing images as single-frame videos to test the generalization of the proposed metric.
+We further extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF Hodosh et al.
+(2013), by viewing images as single-frame videos to test the generalization of the proposed metric.
 The former contains 17K image-caption pairs rated by three human experts on a scale of 1 to 4, while the latter collects binary quality assessments for 48K image-caption pairs from crowd sources.
+Video preprocessing.
+Each video is uniformly sampled into 32 frames and resized to a resolution of $224\times 224$ for both training and testing. For any image dataset, we treat them as one-frame videos.
 Baselines.
-The baselines are organized into three categories: i) Language-based metrics: The evaluation in this category solely relies on the reference texts without considering the visual input.
-Representative metrics are BLEU [26], ROUGE [19], METEOR [3], CIDEr [33], BERTScore [39], SPICE [1], SPICE-Factual [18], Soft-SPICE  [18] and CLAIR [6].
-ii) Image-augmented metrics: These approaches incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore [9], EMScore [31], PAC-S [30], FactVC [21], and G-VEVAL [32].
-They usually support both reference-free and reference-based settings, where the former compares the candidate caption directly to the video without reference captions, while the latter uses both visual and textual references.
+Baselines are organized into three categories: i) Language-based metrics solely rely on text reference without considering the visual input.
+Representative metrics are BLEU Papineni et al.
+(2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), CIDEr Vedantam et al.
+(2015), BERTScore Zhang et al.
+(2020), SPICE Anderson et al.
+(2016), SPICE-Factual Li et al.
+(2023), Soft-SPICE  Li et al.
+(2023) and CLAIR Chan et al.
+(2023b).
+ii) Image-augmented metrics incorporate images as references (alongside reference captions), for better capturing the semantic alignment between the visual input and the candidate caption, e.g., CLIPScore Hessel et al.
+(2022), EMScore Shi et al.
+(2022), PAC-S Sarto et al.
+(2023), FactVC Liu and Wan (2023), and G-VEVAL Tong et al.
+(2025).
+They usually support both reference-free and reference-based settings.
 We report results for both settings for completeness, but our primary focus is on the reference-free setting, which is more practical for real-world, in-the-wild videos.
-iii) Video-based metrics: These methods employ a video encoder to incorporate full video sequences as references.
+iii) Video-based metrics employ a video encoder to incorporate full video sequences.
 To the best of our knowledge, no existing metric in the literature falls into this category.
-Therefore, we adapt CLIPScore [9] to the recent advent of ViCLIP [35] as a stronger baseline, ViCLIPScore.
+Therefore, we adapt CLIPScore Hessel et al.
+(2022) to the recent advent of ViCLIP Wang et al.
+(2024) as a stronger baseline, ViCLIPScore.
 Additionally, we compare VC-Inspector against the base model it builds upon, the vanilla Qwen2.5-VL model, to highlight the benefit of fine-tuning for evaluative reasoning.
+ActivityNet-FG-Eval YouCook2-FG-Eval
+Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
+EMScore Shi et al. (2022) 28.94 40.77 20.21 29.24
+CLIPScore Hessel et al. (2022) 28.10 39.65 18.00 26.14
+Qwen2.5-VL-3B Bai et al. (2025) 37.91 47.80 37.16 47.17
+VC-Inspector-3B 49.53 62.01 44.29 55.31
+Table 1: Correlation scores on the synthetic ActivityNet-FG-Eval Caba Heilbron et al. (2015) and YouCook2-FG-Eval Krishna et al. (2017) datasets. The best score is bolded.
 Metrics.
-In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to human judgments.
-Following prior work [31, 32], we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
+In the caption evaluation literature, the effectiveness of evaluation metrics is typically assessed by measuring their correlation to ground truth.
+Following prior work Shi et al.
+(2022); Tong et al.
+(2025), we report both Kendall’s correlation ($\tau_{b}$) and Spearman’s rank correlation ($\rho$).
 Implementation details.
@@
 During inference, we use a temperature of 0.0 for reproducibility.
+The computational analysis and the training hyperparameters are provided in App.˜B and App.˜C, respectively.
 4.2 Generalization of Quality Estimation
-In this section, we evaluate the consistency of quality estimates using synthetic data, and validate the effectiveness of VC-Inspector by measuring its alignment with human judgments on the video caption evaluation dataset, VATEX-Eval.
-Given the limited availability of video caption datasets, we treat images as single-frame videos and extend our evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF, to assess the generalization.
-Evaluation with the synthetic data.
-ActivityNet-FG-Eval YouCook2-FG-Eval
-Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
-EMScore [31] 28.94 40.77 20.21 29.24
-CLIPScore [9] 28.10 39.65 18.00 26.14
-Qwen2.5-VL-3B [2] 37.91 47.80 37.16 47.17
-VC-Inspector-3B 49.53 62.01 44.29 55.31
-Table 2: Correlation scores on the synthetic ActivityNet-FG-Eval [4] and YouCook2-FG-Eval [14] datasets. The best score is bolded.
-Table 2 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies.
+Evaluation with synthetic data.
+No Reference 1-Reference 9-References
+Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
+Language-based
+BLEU_1 Papineni et al. (2002) - - 12.65 16.52 28.70 36.88
+BLEU_4 Papineni et al. (2002) - - 12.44 36.28 22.76 29.56
+ROUGE-L Lin (2004) - - 12.94 16.89 23.94 31.06
+METEOR Banerjee and Lavie (2005) - - 16.68 21.80 27.64 35.76
+CIDEr Vedantam et al. (2015) - - 17.62 23.02 27.92 36.18
+BERTScore Zhang et al. (2020) - - 15.24 19.82 25.05 32.37
+SPICE Anderson et al. (2016) - - 14.80 18.78 27.41 35.40
+SPICE-factual Li et al. (2023) - - 13.59 17.04 26.05 35.58
+Soft-SPICE Li et al. (2023) - - 21.25 27.61 36.31 46.41
+CLAIR Chan et al. (2023b) - - 36.00 - 34.80 -
+Multimodal - image-based
+CLIPScore Hessel et al. (2022) 22.33 29.09 27.39 35.49 35.21 45.28
+EMScore Shi et al. (2022) 22.88 29.79 28.63 37.05 36.66 47.00
+FactVC Liu and Wan (2023) 22.79 29.69 28.78 37.22 36.18 46.33
+PAC-S Sarto et al. (2023) 25.10 - 32.60 - 31.40 -
+G-VEval Tong et al. (2025) 39.40 - 44.90 - 48.10 -
+Multimodal - video-based
+ViCLIPScore Wang et al. (2024) 30.92 39.86 - - - -
+Qwen2.5-VL-3B Bai et al. (2025) 31.29 36.43 - - - -
+Qwen2.5-VL-7B Bai et al. (2025) 34.70 39.40 - - - -
+VC-Inspector-3B 37.99 42.45 - - - -
+VC-Inspector-7B 42.58 45.99 - - - -
+Table 2: Human correlation scores on the VATEX-EVAL Shi et al.
+(2022) dataset.
+indicates results reported from the original paper.
+The best of each column section is bolded.
+Please note that this work focuses on the No Reference setting.
+Our best model outperforms all other models in this setting, while remaining competitive with reference-based metrics.
+Table˜1 presents the correlation between metric scores and ground truth annotations in the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, both of which contain pseudo captions with varying degrees of factual inaccuracies.
 VC-Inspector, specifically finetuned for factual grounding, consistently outperforms baselines in differentiating incorrect captions.
 Notably, although it is only trained on ActivityNet-FG-It, its quality estimation generalizes effectively across visual domains.
-This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs that fail to converge.
+This suggests that the proposed data generation pipeline produces stable and reliable quality scores, rather than noisy outputs.
+Metric Flickr8K-Expert Flickr8K-CF
+Reference-based
+BLEU_1 Papineni et al. (2002) 32.20 17.90
+BLEU_4 Papineni et al. (2002) 30.60 16.90
+ROUGE Lin (2004) 32.10 19.90
+METEOR Banerjee and Lavie (2005) 41.50 22.20
+CIDEr Vedantam et al. (2015) 43.60 24.60
+SPICE Anderson et al. (2016) 51.70 24.40
+BERTScore Zhang et al. (2020) - 22.80
+CLIPScore Hessel et al. (2022) 52.60 36.40
+PAC-S Sarto et al. (2023) 55.50 37.60
+Reference-free
+CLIPScore Hessel et al. (2022) 51.10 34.40
+PAC-S Sarto et al. (2023) 53.90 36.00
+VC-Inspector-3B 59.86 39.00
+VC-Inspector-7B 63.43 45.97
+Table 3: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF Hodosh et al.
+(2013) dataset.
+The overall best scores are bolded and best of each section is underlined.
+Our model outperforms even the reference-based methods.
 Evaluation on VATEX-Eval.
-Table 5.2 presents our primary results: human correlation scores on the VATEX-EVAL [31] dataset.
+Table 2 presents our primary results: human correlation scores on the VATEX-EVAL Shi et al.
+(2022) dataset.
 Baselines are grouped by reference type as outlined in section 4.1, and the three column-sections correspond to the No Reference, 1-Reference, and 9-Reference settings, respectively.
-Language-based metrics, which rely heavily on textual references, are not applicable to our target No Reference setting.
+Language-based metrics, which rely heavily on textual references, are not applicable to our target, No Reference setting.
 Therefore, we focus our comparisons on multimodal methods that incorporate visual inputs.
-VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on CLIP embeddings, including CLIPScore, EMScore, FactVC, and PAC-S.
-When adapting CLIPScore to the recently introduced ViCLIP [35] (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches.
-Despite ViCLIPScore being a stronger baseline, it still underperforms relatively to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation.
-We further compare VC-Inspector with G-VEval [32], a recent LMM-based evaluation method.
-However, G-VEval is based on GPT-4o, a proprietary, paid model with an estimated 200B parameters.
+VC-Inspector consistently outperforms all evaluated metrics in the reference-free setting, particularly those based on (image-based) CLIP embeddings.
+When adapting CLIPScore to the recently introduced ViCLIP Wang et al.
+(2024) (which adopts a video encoder), we observe a considerable gain over those image-CLIP-based approaches.
+Despite ViCLIPScore being a stronger baseline, it still underperforms relative to VC-Inspector and its underlying model, as the context length of the ViCLIP model (i.e., 32) is much shorter than the 32K context length of ours, which supports a more flexible and comprehensive caption evaluation.
+Furthermore, VC-Inspector outperforms G-VEval (Tong et al., 2025) despite being based on GPT-4o, a substantially larger proprietary model.
+This demonstrates that explicit factual grounding and explanation supervision are more critical than model scale alone.
 In contrast, VC-Inspector is open-source, lightweight, and reproducible, with configurations available at 3B and 7B parameters depending on system requirements.
 While having a compact size, our 7B model achieves the highest correlation to human evaluations and could potentially enable on-the-fly quality estimation during training, making it a viable reward model in Reinforcement Learning (RL) applications.
-Although our primary focus is to compare with the reference-free metrics, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
+Although we focus on the reference-free setting, it is noteworthy that VC-Inspector even surpasses the performance of most reference-based metrics.
 Evaluation on Flickr8K.
-No Reference 1-Reference 9-References
-Metric $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$ $\tau_{b}$ $\rho$
-[…]
-BLEU_1 [26] - - 12.65 16.52 28.70 36.88
-BLEU_4 [26] - - 12.44 36.28 22.76 29.56
-ROUGE-L [19] - - 12.94 16.89 23.94 31.06
-METEOR [3] - - 16.68 21.80 27.64 35.76
-CIDEr [33] - - 17.62 23.02 27.92 36.18
-BERTScore [39] - - 15.24 19.82 25.05 32.37
-SPICE [1] - - 14.80 18.78 27.41 35.40
-SPICE-factual [18] - - 13.59 17.04 26.05 35.58
-Soft-SPICE [18] - - 21.25 27.61 36.31 46.41
-CLAIR [6] - - 36.00 - 34.80 -
-[…]
-CLIPScore [9] 22.33 29.09 27.39 35.49 35.21 45.28
-EMScore [31] 22.88 29.79 28.63 37.05 36.66 47.00
-FactVC [21] 22.79 29.69 28.78 37.22 36.18 46.33
-PAC-S [30] 25.10 - 32.60 - 31.40 -
-G-VEval [32] 39.40 - 44.90 - 48.10 -
-[…]
-ViCLIPScore [9, 35] 30.92 39.86 - - - -
-Qwen2.5-VL-3B [2] 31.29 36.43 - - - -
-Qwen2.5-VL-7B [2] 34.70 39.40 - - - -
-VC-Inspector-3B 37.99 42.45 - - - -
-VC-Inspector-7B 42.58 45.99 - - - -
-Table 3: Human correlation scores on the VATEX-EVAL [31] dataset.
-indicates results reported from [32] due to GPT-4 licensing issue.
-The best score for each column section is bolded.
-Please note that this work focuses on the No Reference setting.
-Our best model has outperformed all other models in this setting, while remaining competitive with metrics that rely on references.
-Metric Flickr8K-Expert Flickr8K-CF
-[…]
-BLEU_1 [26] 32.20 17.90
-BLEU_4 [26] 30.60 16.90
-ROUGE [19] 32.10 19.90
-METEOR [3] 41.50 22.20
-CIDEr [33] 43.60 24.60
-SPICE [1] 51.70 24.40
-BERTScore [39] - 22.80
-CLIPScore [9] 52.60 36.40
-PAC-S [30] 55.50 37.60
-[…]
-CLIPScore [9] 51.10 34.40
-PAC-S [30] 53.90 36.00
-VC-Inspector-3B 59.86 39.00
-VC-Inspector-7B 63.43 45.97
-Table 4: Correlation score ($\tau_{b}$) with human judgments on Flickr8k-Expert and Flickr8k-CF [10] dataset.
-The overall best scores are bolded and best of each section is underlined.
-Our model has outperformed even the reference-based methods.
-To expand our evaluation beyond the limited availability of video caption datasets, we treat images as an extreme case of short videos with a single frame.
-Table 5.2 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF.
+To expand our evaluation beyond the limited availability of video caption datasets, we treat images as single-frame videos.
+Table 3 reports the human correlation score on two widely adopted benchmarks, Flickr8K-Expert and Flickr8K-CF.
 The former requires a more fine-grained differentiation among the captions, with ratings distributed across 4 levels, whereas the latter employs binary judgments.
@@
 This allows the model to perform effectively across both benchmarks, despite their differing rating scheme.
-In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, substantially narrowing the gap to inter-human correlation ($\tau_{b}\approx 73$[1], and even outperforming several reference-based metrics.
+In these evaluations, VC-Inspector remains the best-performing method under the reference-free setting, and even outperforming several reference-based metrics.
 These results demonstrate the strong generalization capability of VC-Inspector across visual domains and video lengths.
-5.3 Ablation and Analysis
-We conduct ablation studies on VC-Inspector components on the VATEX-Eval benchmark:
-Data synthesis strategies.
-Since both objects and actions are informative factual elements for a video, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components.
-To this end, during the data generation process (outlined in section 4.1), we systematically altered both elements in ground truth captions to create pseudo captions for training.
-In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both.
-As shown in Table 5, all variants achieve strong alignment with human ratings compared to EMScore.
-However, the variant that alters both objects and actions yields the best performance.
-These results highlight the importance of both factual elements, object and action, in capturing the context of video content for caption evaluation.
-They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
-Role of explanations.
-Figure 4: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others).
-VC-Inspector produces quality assessments highly consistent with ground truth scores, along with explanatory insights into factual errors (highlighted in red).
+4.3 Ablation on Data Synthesis Strategies
 Metric Data Synthesis $\tau_{b}$ $\rho$
-EMScore [31] n/a 22.88 29.79
+EMScore Shi et al. (2022) n/a 22.88 29.79
 VC-Inspector-3B Change objects only 36.40 41.20
@@
 Change both (Ours) 37.99 42.45
-Table 5: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval.
-During our data generation process, we systematically synthesize pseudo captions along with their corresponding quality estimates.
-This process yields an informative “side product": explanations that indicate where factual errors are located within the candidate captions.
-We leverage these explanations as an auxiliary supervision signal during training, enabling the model to implicitly learn factual grounding, which has been shown to be effective compared to the variant without explanations in Table 6.
-Not only are explanations useful in training, but they also enhance the interpretability of the model-based metric, providing transparency of why the scores are assigned to given captions.
-Further evaluation of the explanation quality is provided in the supplementary.
-Metric stability.
-To evaluate the stability of our metric, following prior work [36], we computed the Pearson correlation between scores obtained across two runs.
-We find a perfect correlation of 1.0 (as expected with temperature 0.0), confirming that VC-Inspector produces stable quality estimation across multiple runs.
+Table 4: Ablation study on synthetic data generation strategies. Human correlation scores ($\tau_{b}$, $\rho$) are reported on the VATEX-Eval benchmark.
+Since both objects and actions are key factual elements in video captions, grounding the evaluator in factual understanding requires instructing the model to identify errors in these components.
+To this end, we systematically altered both elements in ground truth captions during data generation to create pseudo captions for training.
+In this section, we ablate the impact of modifying these elements in the ActivityNet training set by evaluating three variants: i) Changing objects only, ii) Changing actions only, and iii) Changing both.
+As shown in Table 4, all variants exhibit strong alignment with human ratings compared to EMScore Shi et al.
+(2022).
+However, the variant that alters both objects and actions yields the best performance.
+These results underscore the importance of both factual elements, objects and actions, for capturing video context in caption evaluation.
+They also demonstrate the robustness and generalization of the proposed paradigm across different factual errors.
+4.4 Analysis of Explanations
+Impact of explanation on performance.
+Our data generation process produces not only pseudo captions and quality scores but also an informative “side product": explanations identifying factual errors in candidate captions.
+We leverage these explanations as auxiliary supervision during training, enabling the model to learn better factual grounding, which has been shown to be effective compared to the variant without explanations in Table 5.
+Beyond training benefits, explanations enhance the interpretability of the model-based metric, providing transparency of why specific scores are assigned.
+Using explanations for caption refinement.
 VC-Inspector-3B $\tau_{b}$ $\rho$
@@
 Table 5: Impact of explanations on the model performance. In the “Without Explanation" setting, we have trained the model only on the score, removing the pseudo explanations.
-Computational efficiency.
-We assess the computational efficiency by comparing the average runtime per video clip.
-EMScore [31], ViCLIPScore [35], and VC-Inspector require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU.
-VC-Inspector demonstrates superior efficiency compared to existing methods.
-5.4 Qualitative Results
-Figure 4 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (middle, bottom) datasets.
+Having shown that explanations aids in evaluating caption quality, we now ask: can they also help improve the captions themselves?
+To assess this, we adopt Qwen2.5-VL-7B as the captioning model and use VC-Inspector-7B as the critique model that provides feedback during an iterative refinement process, repeated for up to 10 iterations.
+The prompt used for refinement is provided in App.
+D.
+We conduct this experiment with CAPability (Liu et al., 2025), a recent benchmark that evaluates caption quality across 12 distinct dimensions and two modalities (image and video), with human-annotated ground truth.
+Following its protocol, final captions are scored by GPT-4.1.
+Note that we report dimensions related to object semantics and actions in Table˜6, and provide the full results in Table˜8.
+We implement baselines using the same captioner but with different feedback strategies: (a) no feedback, and (b) feedback from the off-the-shelf Qwen2.5-VL-7B.
+Results show that the captioner guided by VC-Inspector consistently improves over both baselines on these dimensions, demonstrating the effectiveness of explanations for caption refinement.
+Notably, comparing the performance of (b) to (a), we find that undesired feedback to the captioner can even hurt the performance, significantly.
+This further underscores the importance of explicit factual grounding in VC-Inspector.
+Evaluation of explanation quality.
+For a more direct assessment of explanation quality, we randomly sample 100 captions from the VATEX-Eval dataset, and evaluate the explanations generated by VC-Inspector-7B, using 3 of the authors as annotators.
+For each explanation, annotators are provided with the corresponding video and candidate caption, and asked to rate the explanation on a Likert scale rather than binary judgments (e.g., good vs.
+bad), as explanations may be correct in some aspects while missing others.
+The explanations achieve a mean score of 4.33 with a standard deviation of 0.8.
+Inter-annotator agreement, measured by Krippendorff’s Alpha (ordinal), is 0.46, indicating moderate agreement.
+Notably, agreement within a margin of one point reaches 85.4%, suggesting strong consistency among annotators.
+In addition to human evaluation, we present language model based evaluation in App.˜B.
+Feedback Provider
+Task None Qwen2.5-VL-7B VC-Inspector-7B
+Dynamic Object Num. 17.4 17.3 20.7
+Action 58.2 18.8 58.6
+Object Category 67.8 58.9 68.0
+Object Number 28.8 20.4 29.3
+Spatial Relation 56.6 45.7 57.1
+Table 6: F1 scores across different tasks. Video-based tasks are highlighted in blue, while image-based tasks are highlighted in beige. Best results are shown in bold.
+5 Conclusion
+This work tackles the challenge of evaluating video captions across diverse domains without relying on human-annotated references.
+We conducted an extensive review of existing metrics, identified their limitations, and proposed VC-Inspector, a novel fact-aware evaluator for reference-free video caption evaluation.
+By incorporating factual analysis into caption assessment, we equipped a lightweight, open-source LMM with the factual grounding capabilities through training on captions of diverse quality, where pseudo captions with controllable factual errors were systematically generated.
+Experimental results across multiple domains demonstrate that VC-Inspector achieves high alignment with human judgments, and outperforms existing metrics in detecting factual errors.
+Its versatility and interpretability make it a practical tool for assessing factual accuracy in real-world video captioning scenarios.
+Furthermore, ablation studies on predicted explanations underscore its effectiveness in quality estimation and caption refinement.
+Limitations
+VC-Inspector primarily targets object and action correctness, which constitute major sources of factual errors in video captions.
+While these factors cover many common failure cases, other important aspects, e.g., attributes, spatial relationships, and fine-grained temporal ordering, are not explicitly modeled and remain directions for future work.
+Additionally, the training process partially relies on synthetically generated captions and pseudo-scores.
+Although these scores are deterministic and validated through correlation with human judgments, they may not fully capture the diversity of real-world captioning errors.
+Finally, while VC-Inspector is substantially more lightweight than proprietary LLM-based evaluators, it still depends on a multimodal backbone, resulting in higher computational costs compared to purely embedding-based metrics.
+Despite these limitations, VC-Inspector represents a meaningful step toward explanation-aware, factually grounded, and reference-free video caption evaluation, providing a flexible foundation for future extensions.
+Appendix
+Appendix A Qualitative Results
+Fig.˜5 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (others) datasets.
 Without relying on reference captions, the model evaluates candidate captions based on the associated video content and produces quality scores that closely mirror human judgments.
-The explanation pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns factually grounded scores accordingly.
-Additional qualitative results are provided in the supplementary.
-6 Conclusion and Discussion
-This work addresses the challenge of evaluating video captions across diverse domains without relying on human-annotated reference captions.
-We identified the limitations of existing metrics, and proposed a novel reference-free and factually grounded evaluation framework, VC-Inspector, based on LMMs.
-By delving into the factual analysis of captions, we empower the LMM with the factual grounding ability through instruction tuning, where we systematically create pseudo captions of diverse quality using an LLM.
-Experimental results across multiple domains demonstrate that VC-Inspector achieves high consistency with human evaluators, and outperforms existing metrics on detecting factual errors.
-Its versatility and interpretability make it a practical tool for evaluating the factual accuracy of video captions in real-world settings.
-While our current focus is on factual errors, particularly object and action substitutions, the proposed paradigm can be potentially extended to simulate other types of errors (e.g., attribute replacements), as supported by our ablations on synthesizing only one type of error.
-Additionally, we recognize the importance of evaluating temporal coherence and narrative structure, which remain underexplored.
-Future work should expand this paradigm to simulate and detect a broader range of errors, enabling more comprehensive and context-aware evaluation strategies for video captions.
-Appendix
-The appendix is organized as follows.
-Section A lists the detailed prompts used in the proposed data generation pipeline, model fine-tuning, and explanation evaluation.
-Section B provides the evaluation of explanations generated by VC-Inspector, and section C presents additional visual examples with candidate captions of diverse quality.
-Appendix A Prompts
+The explanation effectively pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns scores accordingly.
+In Fig.˜6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual errors.
+The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual inaccuracies.
+Appendix B Additional Results
+B.1 Computational Efficiency
+We assess computational efficiency by comparing the average runtime per video clip.
+EMScore (Shi et al., 2022), ViCLIPScore (Wang et al., 2024), and VC-Inspector-3B require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU.
+These results indicate that VC-Inspector achieves superior efficiency compared to existing methods.
+For EMScore and ViCLIPScore, we use the official implementations provided by the authors, while for VC-Inspector we employ vLLM (Kwon et al., 2023) with online serving.
+Computational efficiency is measured over 2,000 videos, each uniformly sampled into 16 frames.
+Figure 5: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others).
+VC-Inspector produces quality assessments consistent with ground truth scores, and explanatory insights into factual errors (highlighted in red).
+Figure 6: Additional visualization of VC-Inspector results on ActivityNet-FG-Eval videos, with candidate captions of diverse quality.
+Incorrect objects and actions are identified by VC-Inspector and labeled in red.
+B.2 Evaluation of Explanation Quality using Language Models
+We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations.
+Table 7 presents two metrics for this experiment.
+The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings.
+On the other hand, the LLM Score is derived by using the prompt detailed in section D to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to Liu et al.
+(2023a).
+We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations.
+Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
+B.3 Complete Results for Caption Refinement
+In §˜4.4, we present the experiment of caption refinement with CAPability benchmark, where we report the dimensions concerning objects and actions, as they are the primary focus of this work.
+We provide the F1 scores of all the 12 dimensions in Table˜8 for completeness.
+Dataset BERT Score LLM Score
+ActivityNet-FG-Eval 0.79 93.11
+YouCook2-FG-Eval 0.70 90.97
+Table 7: Evaluation of the explanations generated by VC-Inspector-3B on two synthetic evaluation datasets.
+Feedback Provider
+Task None Qwen2.5-VL-7B VC-Inspector-7B
+Camera Angle 37.5 6.4 36.1
+Camera Movement 26.4 0.6 29.3
+Character Identification 29.6 17.3 30.5
+Dynamic Object Num. 17.4 17.3 20.7
+Event 58.2 18.8 58.6
+Object Category 67.8 58.9 68.0
+Object Color 73.0 51.9 72.3
+Object Number 28.8 20.4 29.3
+OCR 69.6 62.9 68.5
+Scene 63.5 63.0 63.2
+Spatial Relation 56.6 45.7 57.1
+Style 84.3 78.7 84.2
+Average 51.1 36.8 51.5
+Table 8: F1 scores across different tasks. Video-based tasks are highlighted in blue, while image-based tasks are highlighted in beige. Best results are shown in bold.
+Appendix C Training Hyperparameters
+We report all training hyperparameters in Table˜9.
+Hyperparameter Value
+Training Configuration
+Epochs 1
+Global Batch Size 128
+Learning Rate 1e-4
+LR Scheduler Cosine (min: 1e-5)
+Warmup Ratio 0.05
+LoRA Configuration
+Rank ($r$) 32
+Alpha ($\alpha$) 32
+Dropout 0.05
+Module Freezing
+Vision Encoder
+Language Model
+Projector
+Video Processing
+Number of Frames 32
+Max. Pixels $224\times 224$
+Table 9: Hyperparameters for instruction fine-tuning.
+Appendix D Prompts
 The data generation prompts are reported on the following blocks:
-  • Extract object – Prompt A
-  • Extract action – Prompt A
-  • Find similar object – Prompt A
-  • Find similar action – Prompt A
-  • Substitute object or action – Prompt A
-The fine-tuning prompt and the prompt for evaluating the generated explanations are reported in Prompt A and Prompt A, respectively.
-Prompt A.1: Prompt to extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", …].
+  • Extract object – Prompt D
+  • Extract action – Prompt D
+  • Find similar object – Prompt D
+  • Find similar action – Prompt D
+  • Substitute object or action – Prompt D
+The data generation process provides the information of incorrect objects and actions. We format the explanations of these factual errors using the following template:
+  • Captions without errors: The caption is helpful, relevant, accurate, and informative to the video content.
+  • Captions with errors: The caption does not accurately capture the video content.
+For example, the actions ({{wrong_act}}) are incorrect / the objects ({{wrong_obj}}) are incorrect / the objects ({{wrong_obj}}) and actions ({{wrong_act}}) are incorrect.
+These explanations are then used in Prompt D for training VC-Inspector.
+The prompt for evaluating the generated explanations is in Prompt D. Prompt D is used to refine the caption based on the explanation.
+Prompt D.1: Extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", ...].
 Don’t include any verbs.
 ONLY REPLY THE ANSWER.
-### Input: {{caption}}
-### Output
-Prompt A.2: Prompt to extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", …]. ONLY REPLY THE ANSWER.
-### Input: {{caption}}
-### Output
-Prompt A.3: Prompt to find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent.
+### Input: {{caption}} ### Output
+Prompt D.2: Extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", ...].
+ONLY REPLY THE ANSWER.
+### Input: {{caption}} ### Output
+Prompt D.3: Find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent.
 The new class cannot be a synonym or similar terms to the original object.
@@
 ONLY REPLY THE NEW CLASS.
-### Input: {{object}}
-### Output:
-Prompt A.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action.
+### Input: {{object}} ### Output:
+Prompt D.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action.
 The new action cannot be a synonym or similar terms to the original action.
@@
 ONLY REPLY THE NEW ACTION.
-### Input: {{action}}
-### Output:
-Prompt A.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act}} in {{caption}} as {{new_obj_act}} .
+### Input: {{action}} ### Output:
+Prompt D.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act} in {{caption}} as {{new_obj_act}}.
 Keep the answer in the same format as {{caption}}.
 ONLY REPLY THE ANSWER.
-### Input: cap
-### Output:
-Prompt A.6: Fine-tuning prompt ### USER:
-[[VIDEO]]
-<caption> {{caption}} </caption> You are given a video and a caption describing the video content.
+### Input: {{cap}} ### Output:
+Prompt D.6: Fine-tuning prompt ### USER: {{VIDEO}} <caption>{{caption}}</caption> You are given a video and a caption describing the video content.
 Please rate the helpfulness, relevance, accuracy, level of details of the caption.
@@
 STRICTLY FOLLOW THE FORMAT.
-### ASSISTANT:
-{{quality_score}}
-The caption does not accurately capture the video content.
-For example, the actions ( {{wrong_act}} ) are incorrect / the objects ( {{wrong_obj}} ) are incorrect/the objects ( {{wrong_obj}} ) and actions ( {{wrong_act}} ) are incorrect.
-Prompt A.7: Prompt for explanation evaluation [Context]
-{{ground_truth_caption}}
-[Caption]
-{{caption_to_evaluate}}
-[Groundtruth]
-{{ground_truth_explanation}}
-[End of Groundtruth]
-[Assistant]
-{{predicted_explanation}}
-[End of Assistant]
-[System]
-We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video.
+### ASSISTANT: {{quality_score}} {{explanation}}
+Prompt D.7: Prompt for explanation evaluation [Context] {{ground_truth_caption}} [Caption] {{caption_to_evaluate}} [Groundtruth] {{ground_truth_explanation}} [End of Groundtruth] [Assistant] {{predicted_explanation}} [End of Assistant] [System] We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video.
 For your reference, the visual content in the video is represented with a few sentences describing the same video.
@@
 Provide an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
-Please first output a single line containing the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
-Appendix B Evaluation on Explanation Quality
-We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations.
-Table 7 presents two metrics for this experiment.
-The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings.
-On the other hand, the LLM Score is derived by using the prompt detailed in section A to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to [20].
-We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations.
-Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
-Dataset BERT Score LLM Score
-ActivityNet-FG-Eval 0.79 93.11
-YouCook2-FG-Eval 0.70 90.97
-Table 7: Evaluation of the explanations generated by VC-Inspector-3B on two synthetic evaluation datasets.
-Figure 5: Additional visual examples on VATEX-Eval. VC-Inspector identifies the incorrect objects (highlighted in red) and assigns scores aligned with human evaluators.
-Figure 6: Additional visualization of VC-Inspector results on ActivityNet-FG-Eval videos, with candidate captions of diverse quality.
-Incorrect objects and actions are identified by VC-Inspector and labeled in red.
-Appendix C Additional Visual Examples
-Figure 5 presents additional visual examples of VC-Inspector applied to VATEX-Eval.
-Consistent with the results shown in Figure 4 of the main paper, VC-Inspector effectively identifies incorrect objects and actions in candidate captions, assigning quality scores that closely align with human judgments.
-In Figure 6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual inaccuracies.
-The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual errors.
+Please first output a single line containing the score.
+In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
+Prompt D.8: Prompt to refine caption based on explanation Previous Caption: {{caption}} Evaluation: {{explanation}} Based on the previous caption and evaluation, refine the caption to make it more accurate and detailed.
+Use the evaluation to find out the missing details and use the video frames to fix the missing details.
+Just output the refined caption, without any other text.
+Appendix E Use of AI Assistance
+During the preparation of this work, the authors utilized Cursor [^2] https://cursor.com for coding assistance and ChatGPT [^3] https://chatgpt.com/ for proofreading and language refinement.
+All input provided to these tools consisted of original content authored by the authors.